Determine the specific heat of lead if :

a- A 150 g sample of lead is heated to the temperature of boiling water (100 oC).

b- A 50 g sample of water is added to a thermally isolated beaker, and its temperature is found to be 20 oC.

c- The hot lead is dumped into cold water, and the final lead-water mixture has a temperature of 28.8 oC.

1
Expert's answer
2021-06-14T13:37:01-0400

Let c be the specific heat of Pb


Heat lost by Pb = mass x specific heat x temperature change of Pb


= 100.0 x c x (100.0 - 26.4) = 7360c J



Heat gained by water = mass x specific heat x temperature change of water


= 150 x 4.184 x (26.4 - 25.0) = 878.64 J



Total heat lost = total heat gained


7360c = 878.64



Specific heat of capacity of Pb = c = 878.64/7360


= 0.119 J/g C



Molar heat capacity of Pb = specific heat capacity x molar mass of Pb


= 0.119 x 207.2


= 24.7 J/mol C